Dave currently consumes 10 hotdogs and 6 sodas each week. At his current consumption basket , his marginal utility for hot dogs is 5 and his marginal utility for soda is 3. if the price of one hot dog is $1 and the price of one soda is $0.50, is Dave currently maximizing his utility? if not, how should he reallocate his spending in order to increase his utility?
Answer #1

At the utility maximization bundle : marginal utility per dollar are equal for each good.

MUD/ PD = 5/1 = 5

MUS/PS =3/0.50 = 6

No , Dave is currently not maximizing his utility .

To maximize his utility , Dave should consume more sodas and fewer hot dogs. As a result of this , marginal utility for sodas decreases and marginal utility for hot dogs increases. And Dave maximize utility at which marginal utility per dollar is equal for both goods.